# DUMP: Automated Distribution-Level Curriculum Learning for RL-based LLM Post-training

### v2

@@ 2 Background
$\mathcal{J}_{\text{GRPO}}(\theta)=\mathbb{E}_{x,\{o_{i}\}}\left[\frac{1}{G}\sum_{i=1}^{G}\min\left(\frac{\pi_{\theta}(o_{i}|x)}{\pi_{\text{ref}}(o_{i}|x)}\hat{A}_{i},\text{clip}\left(\frac{\pi_{\theta}(o_{i}|x)}{\pi_{\text{ref}}(o_{i}|x)},1-\epsilon,1+\epsilon\right)\hat{A}_{i}\right)-\beta\,\mathbb{D}_{\text{KL}}(\pi_{\theta}\|\pi_{\text{ref}})\right],$ (2)
where $\pi_{\theta}(o_{i}|x)$ is the probability assigned by the current model to output $o_{i}$, and $\pi_{\text{ref}}(o_{i}|x)$ is the same under the reference model. The first term inside the summation is a clipped policy ratio scaled by $\hat{A}_{i}$, similar to PPO [23], which prevents overly large updates. The outer expectation is taken over prompts $x$ and their sampled output groups $\{o_{i}\}$. The second term is a KL divergence penalty that regularizes the updated policy $\pi_{\theta}$ to stay close to $\pi_{\text{ref}}$, weighted by a hyperparameter $\beta$. This formulation eliminates the need for an explicit value baseline and stabilizes training by comparing outputs within local groups.
@@ 3.2 Formalizing Distribution-Level Curriculum Learning as Multi-armed Bandit
4:         $\mathcal{A}_{d_{j}}^{w}\leftarrow\text{[ ]}$triangleright Sliding window for absolute advantages
5:         $n_{d_{j}}\leftarrow 0$triangleright Total samples seen from $d_{j}$
6:         $P(d_{j})\leftarrow\frac{1}{N}$triangleright Equal initial weights
@@
11:              $n_{d_{j}}\leftarrow n_{d_{j}}+|\mathcal{B}_{t,d_{j}}|$triangleright Update sample count; $\mathcal{B}_{t,d_{j}}$: subset of batch from $d_{j}$
12:              $\mathcal{A}_{d_{j}}^{w}\leftarrow\mathcal{A}_{d_{j}}^{w}\cup\left\{|\hat{A}(o)|\mid x\in\mathcal{B}_{t,d_{j}},\,o\sim\pi_{\theta}(\cdot|x)\right\}$triangleright Append new advantages from $d_{j}$
13:              $\mathcal{A}_{d_{j}}^{w}\leftarrow\mathcal{A}_{d_{j}}^{w}[-k:]$triangleright $k$: Window Size; Keep last $k$ elements
@@
17:              $\hat{L}(d_{j})\leftarrow\frac{1}{|\mathcal{A}_{d_{j}}^{w}|}\sum_{a\in\mathcal{A}_{d_{j}}^{w}}a$triangleright Mean of absolute advantages
18:              $\text{UCB}(d_{j})\leftarrow\hat{L}(d_{j})+\sqrt{\frac{2\log(n_{\text{total}}+1)}{n_{d_{j}}+1}}$triangleright Eq. 3
@@
20:         $P(d_{j})\leftarrow\frac{\exp(\text{UCB}(d_{j})/\tau)}{\sum_{j=1}^{N}\exp(\text{UCB}(d_{j})/\tau)}\quad\forall d_{j}\in\mathcal{D}$triangleright $\tau$: temperature
@@ 4.1 Experiments Setup
  • Format Reward. A binary reward based on whether the output strictly adheres to the expected format. If the model includes exactly one well-formed <think> and one <answer> section in the correct order, it receives a reward of $+1$; otherwise, it receives a penalty of $-1$.
  • Answer Reward. We evaluate the correctness of the final answer. If the predicted identities fully match the ground truth, the model receives a reward of $+2$; if the answer is incorrect, $-1.5$; and if the answer is missing or unparsable, $-2$.
@@
Other Implementation Details. All experiments are conducted on servers equipped with 8 Nvidia A100 GPUs. Our method is implemented with VeRL [41] LLM Reinforcement Learning framework. We use GRPO [3] as the training algorithm and follow standard practice for actor rollout and optimization. The actor learning rate is set to $1e{-6}$, training batch size is set to 128, and the PPO mini-batch size is 32. KL divergence regularization is applied to encourage alignment with the reference policy, with a KL loss coefficient of 0.001. Each rollout batch contains 16 responses. If not specified, we allow for a maximum response length of 20480 and 4096 tokens during training for Qwen2.5-7B-Instruct-1M and Qwen2.5-3B-Instruct, respectively. The window size $k$ and the temperature $\tau$ in our curriculum learning framework is set to 300 and 0.1, respectively.
@@ 4.3 Ablation Study on the Sampling Strategy
In this section, we ablate the sampling strategy used in DUMP’s UCB-based scheduler. As described in Algorithm 1, our method applies soft sampling controlled by a temperature parameter. The greedy variant (temperature = 0) always selects the distribution with the highest UCB score, while our default uses a small temperature (0.1) to enable probabilistic sampling. We conduct experiments under Setting 1, with a maximum training response length of 10240 tokens. After 100 training steps, the greedy strategy significantly underperforms due to its lack of exploration—it tends to lock onto a single distribution early and fails to adapt. For instance, on the 13- and 14-character K&K tasks, the greedy variant achieves test answer rewards of $-0.91$ and $-1.38$, while soft sampling reaches $-0.66$ and $-1.16$, respectively. These results highlight the importance of maintaining exploration via a non-zero temperature to prevent the scheduler from collapsing onto suboptimal distributions.

### v3

@@ 2 Background
$\mathcal{J}_{\text{GRPO}}(\theta)=\mathbb{E}_{x,\{o_{i}\}}\left[\frac{1}{G}\sum_{i=1}^{G}\min\left(\frac{\pi_{\theta}(o_{i}|x)}{\pi_{\text{old}}(o_{i}|x)}\hat{A}_{i},\text{clip}\left(\frac{\pi_{\theta}(o_{i}|x)}{\pi_{\text{old}}(o_{i}|x)},1-\epsilon,1+\epsilon\right)\hat{A}_{i}\right)-\beta\,\mathbb{D}_{\text{KL}}(\pi_{\theta}\|\pi_{\text{ref}})\right],$ (2)
where $\pi_{\theta}(o_{i}|x)$ is the probability assigned by the current model to output $o_{i}$, $\pi_{\text{old}}(o_{i}|x)$ is the same under the model from previous step, and $\pi_{\text{ref}}(o_{i}|x)$ is that under the reference model. The first term inside the summation is a clipped policy ratio scaled by $\hat{A}_{i}$, similar to PPO [23], which prevents overly large updates. The outer expectation is taken over prompts $x$ and their sampled output groups $\{o_{i}\}$. The second term is a KL divergence penalty that regularizes the updated policy $\pi_{\theta}$ to stay close to $\pi_{\text{ref}}$, weighted by a hyperparameter $\beta$. This formulation eliminates the need for an explicit value baseline and stabilizes training by comparing outputs within local groups.
@@ 3.2 Formalizing Distribution-Level Curriculum Learning as Multi-armed Bandit
[…]
5:   $n_{d_{j}}\leftarrow 0$ $\triangleright$ Total samples seen from $d_{j}$
6:   $P(d_{j})\leftarrow\frac{1}{N}$ $\triangleright$ Equal initial weights
@@
11:     $n_{d_{j}}\leftarrow n_{d_{j}}+|\mathcal{B}_{t,d_{j}}|$ $\triangleright$ Update sample count; $\mathcal{B}_{t,d_{j}}$: subset of batch from $d_{j}$
12:     $\mathcal{A}_{d_{j}}^{w}\leftarrow\mathcal{A}_{d_{j}}^{w}\cup\left\{|\hat{A}(o)|\mid x\in\mathcal{B}_{t,d_{j}},\,o\sim\pi_{\theta}(\cdot|x)\right\}$ $\triangleright$ Append new advantages from $d_{j}$
13:     $\mathcal{A}_{d_{j}}^{w}\leftarrow\mathcal{A}_{d_{j}}^{w}[-k:]$ $\triangleright$ $k$: Window Size; Keep last $k$ elements
@@
17:     $\hat{L}(d_{j})\leftarrow\frac{1}{|\mathcal{A}_{d_{j}}^{w}|}\sum_{a\in\mathcal{A}_{d_{j}}^{w}}a$ $\triangleright$ Mean of absolute advantages
18:     $\text{UCB}(d_{j})\leftarrow\hat{L}(d_{j})+\sqrt{\frac{2\log(n_{\text{total}}+1)}{n_{d_{j}}+1}}$ $\triangleright$ Eq. 3
@@
20:   $P(d_{j})\leftarrow\frac{\exp(\text{UCB}(d_{j})/\tau)}{\sum_{j=1}^{N}\exp(\text{UCB}(d_{j})/\tau)}\quad\forall d_{j}\in\mathcal{D}$ $\triangleright$ $\tau$: temperature
@@ 4.1 Experiments Setup
  • Format Reward. A binary reward based on whether the output strictly adheres to the expected format. If the model includes exactly one well-formed <think> and one <answer> section in the correct order, it receives a reward of +1+1; otherwise, it receives a penalty of 1-1.
  • Answer Reward. We evaluate the correctness of the final answer. If the predicted identities fully match the ground truth, the model receives a reward of +2+2; if the answer is incorrect, 1.5-1.5; and if the answer is missing or unparsable, 2-2.
@@
Other Implementation Details. All experiments are conducted on servers equipped with 8 Nvidia A100 GPUs. Our method is implemented with VeRL [41] LLM Reinforcement Learning framework. We use GRPO [3] as the training algorithm and follow standard practice for actor rollout and optimization. The actor learning rate is set to 1e61e{-6}, training batch size is set to 128, and the PPO mini-batch size is 32. KL divergence regularization is applied to encourage alignment with the reference policy, with a KL loss coefficient of 0.001. Each rollout batch contains 16 responses. If not specified, we allow for a maximum response length of 20480 and 4096 tokens during training for Qwen2.5-7B-Instruct-1M and Qwen2.5-3B-Instruct, respectively. The window size $k$ and the temperature $\tau$ in our curriculum learning framework is set to 300 and 0.1, respectively.
@@ 4.3 Ablation Study on the Sampling Strategy
In this section, we ablate the sampling strategy used in DUMP’s UCB-based scheduler. As described in Algorithm 1, our method applies soft sampling controlled by a temperature parameter. The greedy variant (temperature = 0) always selects the distribution with the highest UCB score, while our default uses a small temperature (0.1) to enable probabilistic sampling. We conduct experiments under Setting 1, with a maximum training response length of 10240 tokens. After 100 training steps, the greedy strategy significantly underperforms due to its lack of exploration—it tends to lock onto a single distribution early and fails to adapt. For instance, on the 13- and 14-character K&K tasks, the greedy variant achieves test answer rewards of 0.91-0.91 and 1.38-1.38, while soft sampling reaches 0.66-0.66 and 1.16-1.16, respectively. These results highlight the importance of maintaining exploration via a non-zero temperature to prevent the scheduler from collapsing onto suboptimal distributions.